OpenStudy (anonymous):

PLEASE HELP FAST!! Dogs in the GoodDog Obedience School win a blue ribbon for learning how to sit, a green ribbon for learning how to roll over, and a white ribbon for learning how to stay. There are 100 dogs in the school. 73 have blue ribbons, 39 have green ribbons, and 62 have white ribbons. 21 have a blue ribbon and a green ribbon; 28 have a green ribbon and a white ribbon; 41 have a blue ribbon and a white ribbon. 14 have all three ribbons. b) How many dogs have not learned any tricks?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

The answer is 2. From the set theory, \[n(A \cup B \cup C) = n(A) + n(B) + n(C) - n(A.B)-n(B.C)-n(C.A) + n(A.B.C)\] The dot implies intersection. You need to add x to the R.H.S for dogs which did not learn any tricks.

OpenStudy (amistre64):

or a venn worked from the inside out will be a good chart to fill in

OpenStudy (anonymous):

Yeah venn diagrams are good. But they are very difficult to draw here.

OpenStudy (amistre64):

Dogs in the GoodDog Obedience School win a blue ribbon for learning how to sit, a green ribbon for learning how to roll over, and a white ribbon for learning how to stay. There are 100 dogs in the school. 73 have blue ribbons, 39 have green ribbons, and 62 have white ribbons. 14 bgw 21-14 bg = 7 28-14 gw = 14 41-14 bw = 27 73 have blue ribbons 39 have green ribbons 62 have white ribbons. 

OpenStudy (amistre64):

adding up all the parts and subtracting the total number of participants leaves how many are not ribbon winners

OpenStudy (anonymous):

Yeah,add all the parts and subtract from total number. correct.

OpenStudy (amistre64):

if we try to reconstruct the original solution process: (bgw) + (bgw-bg) + (bgw-bw) + (bgw-gw) + (b-bgw - bw - bg) + (g-bgw - gw - bg) + (w-bgw - bw - gw) (bgw) + (bgw) -(bg+bw+gw) + (b+g+w) - (3bgw +2bw +2gw +2bg) (2bgw) -(bg+bw+gw) + (b+g+w) - (3bgw +2bw +2gw +2bg) (bgw) + (b+g+w) - 3(bw +gw +bg) now to determine if ive messed up someplace lol

OpenStudy (amistre64):

... personally, id stick with the venns for a set of 3 or less :)

OpenStudy (anonymous):

the venn diagram probably is the simplest way :) label the unknown area as x and that is probably it

OpenStudy (amistre64):

with a little thought to it: the parts remaining are: b - bg - bw + bgw; since bgw is actually subtracted twice 14 bgw 21-14 bg = 7 28-14 gw = 14 41-14 bw = 27 73 - 21 - 41 + 14 = 25 39 - 21 - 28 + 14 = 4 62 - 28 - 41 + 14 = 7 100 - (14+7+14+27+25+4+7) = 100 - 98